OpenStudy (anonymous):

Find the sum of all multiples of 7 between 400 and 500.

OpenStudy (kainui):

Any ideas?

OpenStudy (anonymous):

Well, I'm stuck myself and that is all the information that was provided. Other than that, the answer at the back of the book is 6,321.

OpenStudy (anonymous):

do you know the formula?

OpenStudy (kainui):

There is no need to have formulas for this.

OpenStudy (anonymous):

the its hard

OpenStudy (kainui):

Well you can use a formula, but you won't learn anything other than how to replace letters with numbers.

OpenStudy (anonymous):

the equation should be 

OpenStudy (anonymous):

Sn = n÷2 (a+l) or Sn = n÷2 {2a+(n-1)d}

OpenStudy (anonymous):

Those two are the equations. I used both but I'm pretty sure I made some sort of a mistake

OpenStudy (anonymous):

well then what would have been your first step??

hartnn (hartnn):

which will be the very first number, after 400, which is a multiple of 7 ? can you find that ?

OpenStudy (anonymous):

406

hartnn (hartnn):

right, and which will be the last number? just before 500, which is divisible by 7 ?

OpenStudy (anonymous):

it's 497

hartnn (hartnn):

correct, so we have our series from 406,413,420,... ..497 right ? can you tell, how may terms will be there in all ?

OpenStudy (anonymous):

58 × 7 = 406 71 × 7 = 497 then find the mean. (406 + 497)/2 = 451.5 well there is 17 terms in this sequence so 451.5 * 17 = 6321

hartnn (hartnn):

he/she used the formula : \(\Large S_n = (n/2) [a_1 +a_n]\)

hartnn (hartnn):

which is sum for n terms in an arithmetic series

OpenStudy (anonymous):

Oh! I got it :O

hartnn (hartnn):

glad to hear that :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

But, how do I find out how many terms there are? Oh no. -.- That was a dumb question. Got it.

hartnn (hartnn):

i get 14 terms, how did u get 17 terms ?

OpenStudy (anonymous):

no your right 14 i messed up on my paper i got mixed up

hartnn (hartnn):

It was a good question, i find it like this : range = 500-400 = 100 100/7 = 14 (rounded) alternate way : 497 is the n'th term, common difference = 7 , a1 = 1st term =406 497 = 406 + 7(n-1) n=14

hartnn (hartnn):

i used the n'th term formula there \(a_n = a_1 +(n-1)d\)

OpenStudy (anonymous):

Thanks both of you :)

hartnn (hartnn):

welcome ^_^

OpenStudy (goformit100):

▬ ▬ Sir/Ma'am, A Warm Welcome To Open Study ▬ ▬